A concrete driveway measures 18 yards by 4 yards. The driveway will be 0.5 yards thick. (a) How many cubic yards of concrete will be needed? (b) How much will it cost if concrete is $3.50 per cubic yard?

Answer by jojo14344(1513)

There you go, if you want 0.5 yard thick of concrete on the driveway too (just like the thickness of the driveway), multiply Length, Width, & Thickness = 18*4*0.5=36 cubic yards ------ This how much concrete you need.
And, $3.50/cu.yard * 36 cub.yards = $126 -------- cost for the concrete for the whole driveway (plus labor$$$).
Thank you,
Jojo